OpenStudy (anonymous):

8t^4/40t simply by removing a factor equal to 1

OpenStudy (callisto):

\[\frac{8t^4}{40t} = \frac{8t^3 \times t}{8 \times 5t} \] Can you cancel the common factors in both numerator and denominator?

OpenStudy (anonymous):

so you would cancel the 8 on top and bottom and the t?